Calculate ΔH at 400oC for the reaction

SO2(g) + 1/2O2(g)⟶SO3(g)

given that the standard heat formation at 25oC of SO2 and SO3 are -70.9 and -94.4 (in kcal mol-1) repectively and that the heat capacities in JK-1mol-1 are:

SO2: Cp = 10.39 + 2.54 x 10-3T -1.42 x 10-5T2

SO3: Cp = 13.70 + 6.42 x 10-3T -3.12 x 10-5T2

O2: Cp = 7.16 + 1.00 x 10-3T -0.40 x 10-5T2

1
Expert's answer
2021-06-21T05:28:42-0400

formation looks like this

ΔHrxn=∑(n⋅ΔH0f products)−∑(m⋅ΔH0f reactants)

n and m are the stoichiometric coefficients (number of moles) of the products and reactants, respectively.

ΔH0f SO2=-296.84 kJ/mol

ΔH0f SO3=-395.7 kJ/mol

ΔH0f O2=0 kJ/mol

However this enthalpy change involves Cp and SHC have been provided.

Therefore, the enthalpy change of the reaction will be

∆H°T = ∆)H°T1 + "\intop"T ∆Cp° (T1)dT1-

400-25 = 375°

Cp SO2 + O2 = 12.93 ×10-3T -1.42 ×10-5T2 + 20.12 ×10-3T- 0.4 × 10-5T2

= 17.55 + 3.4(10-3 T) - 1.82(10-5) T2

Cp = a + bT + CT-2 + dT2

Cp = -3.85 + 3.02(10-3)T - 1.3(10-5)T2

Cp = -3.85 + 0.00303T -0.000013T2

= -3.85 + 0.00303(-70.9) -(0.000013(-94.4)

= -3.85 - 0.0215 + 0.00123

Cp = -3.87

∆H = 400 + (-3.87(-395.7 +296.84)

= 400 + -381.195

= 18.805 kJ